3. Why do He and H produce different line spectra? Is the same fundamental process is responsible for emission of photons in both elements? 4. When observing the various metal salts, you probably noticed that any red bands always appear at approximately the same scale reading. The same is true for other colors too. Why?
Answer #1

3-

Theformula for calculating wavelengths is given by Rydberg equation

1/wavelength = Z^2 x Rh ( 1/n1^2 - 1/n2^2)

where n1 = lower energy level,   n2= higher energy level . Emission happens from n2 to n1 level.

Rh is Rydberg constant ,Z is atomic number of atom

H has Z value 1 while He has Z value 2 . Thus based on value of Z we see that our wavelength of emission changes for H and He for same values of n1 , n2 .

The reason is that energy of electron in same orbital changes for different atoms. This is due to change in nuclear charge. H has 1 proton while He has 2 protons . Thus nuclear attraction over electrons is higher in case of He than in H atom.

yes same funda proces is respo for emissions of photons iin both elements.

E=h*f=E_2-E_1